# MultiConIR: Towards multi-condition Information Retrieval

### v2

@@ D.2 Complete Results Of Task 2
D.3 Complete Results Of Document Length
@@
Appendix E Findings In Constructing MultiCon Dataset
E.1 The Use Of LLM-generated Data in Retrieval
In recent years, artificial datasets generated by LLMs have become a common practice for training and evaluating retrieval models (Su et al., 2024; Lee et al., 2024; Weller et al., 2024a). For instance, E5-Mistral (Wang et al., 2024a) rely entirely on LLM-generated datasets for fine-tuning. While this approach can significantly expand training corpora, prior studies have highlighted its potential drawbacks, including introducing inherit linguistic biases of the underlying LLMs (Shumailov et al., 2024), potentially constraining the retrieval model’s performance and generalizability. Furthermore, purely artificial data often lacks the contextual richness and complexity found in real-world retrieval scenarios (Li et al., 2023c; Wang et al., 2024b), making it difficult to capture the actual needs of users’ queries accurately.
@@ E.1 The Use Of LLM-generated Data in Retrieval
E.2 Impact of Different Hard Negative Construction Strategies

### v3

@@ F.1 Complete Results Of Task 2
F.2 Complete Results Of Document Length
@@
F.3 Comparative Results for other LLMs
We evaluated two additional LLMs—Qwen3-32B and Gemini-2.0-Flash—on the same MultiConIR Task 1. As shown in Table 16, the results indicate that Gemini-2.0-Flash attains the highest WinRate (91.20), marginally outperforming GPT-4o (90.47), with Qwen3-32B in third place (86.61). GPT-4o is therefore not significantly superior, confirming that MultiConIR does not inherently favor GPT-4o.
Model Q1 Q2 Q3 Q4 Q5 Q6 Q7 Q8 Q9 Q10 Avg.
Qwen3-32B 91.20 90.64 89.55 88.08 86.91 85.35 84.74 84.44 83.27 81.88 86.61
GPT-4o 95.49 94.89 93.71 92.11 90.81 89.08 88.43 88.08 86.82 85.26 90.47
Gemini-2.0-flash 95.67 95.10 94.06 92.63 91.50 89.99 89.38 89.08 87.95 86.61 91.20
Table 16: Performance comparison across Q1–Q10 queries and average score.
Appendix G Findings In Constructing Multiconir Dataset
G.1 The Use Of LLM-generated Data in Retrieval
In recent years, artificial datasets generated by LLMs have become a common practice for training and evaluating retrieval models (Weller et al., 2024a; Shi et al., 2025, 2024; Zhang et al., 2025; Han et al., 2025). For instance, E5-Mistral (Wang et al., 2024a) rely entirely on LLM-generated datasets for fine-tuning. While this approach can significantly expand training corpora, prior studies have highlighted its potential drawbacks, including introducing inherit linguistic biases of the underlying LLMs (Shumailov et al., 2024), potentially constraining the retrieval model’s performance and generalizability. Furthermore, purely artificial data often lacks the contextual richness and complexity found in real-world retrieval scenarios (Li et al., 2023c; Wang et al., 2024c), making it difficult to capture the actual needs of users’ queries accurately.
@@ G.1 The Use Of LLM-generated Data in Retrieval
G.2 Impact of Different Hard Negative Construction Strategies